OpenStudy (anonymous):

For the equation 16x^2 + 25y^2 -96x - 200y = -144, how do I write this in standard form? This is an ellipse.

OpenStudy (anonymous):

@jim_thompson5910 @Directrix @Nnesha @perl

OpenStudy (anonymous):

i got stuck and got 16(x-3)^2 + 25(y-2)^2 = 100. What do I do next?

OpenStudy (anonymous):

@Hero @e.mccormick @freckles

OpenStudy (freckles):

first isn't 200/25=8 ? \[16(x^2-6x)+25(y^2-?y)=-144 \\ 16(x^2-6x+9)+25(y^2-?y+?)=-144+16(9)+25(?) \\ \]

OpenStudy (anonymous):

oh yeah, lolz my bad thnx @freckles

OpenStudy (e.mccormick):

It will end up looking somethng like: \(\dfrac{(x+h)^2}{a^2}+\dfrac{(y+k)^2}{b^2}=1\)

OpenStudy (freckles):

\[16(x^2-6x)+25(y^2-8y)=-144 \\ 16(x^2-6x+9)+25(y^2-8y+?)=-144+16(9)+25(?)\] so what would that one question mark need to be to complete the square for the y part

OpenStudy (anonymous):

16

OpenStudy (freckles):

so you add 25(16) on one side so we need to do that to the other actually already put that one 25( ) over there just left that other spot blank for the 16

OpenStudy (freckles):

\[16(x^2-6x+9)+25(y^2-8y+16)=-144+16(9)+25(16)\]

OpenStudy (anonymous):

ok i got ((x-3)^2/25) + ((y-2)^2/16)

OpenStudy (anonymous):

=1

OpenStudy (freckles):

but that 2 should be 4 ?

OpenStudy (anonymous):

oh yeah forgot to fix that

OpenStudy (anonymous):

thnx @freckles